OpenStudy (anonymous):

I haven't done Algebra in over 5 years, so I need help with this problem: What’s the equation of a line that passes through points (0, -1) and (2, 3)?

OpenStudy (anonymous):

First to find the slope of the line you use the equation: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] pick one of the pairs to plug into \[y _{2}\] and \[x _{2}\] just make sure to keep the order after you start plugging numbers in. Does this make sense so far?

OpenStudy (anonymous):

you want to obtain y=mx+b and that formula is the formula for slope (aka "m") @masukasu

OpenStudy (anonymous):

How do I know what numbers to assign to y and x?

OpenStudy (anonymous):

so this is what it will look like once you plug the points in: \[\frac{ (3) - (-1) }{ (2)-(0) }\]

OpenStudy (anonymous):

You have the points (0, -1) and (2, 3); you know that 0 and 2 are x-values and -1 and 3 are y-values, correct?

OpenStudy (anonymous):

I used the second set of points to use first only so the equation won't look messy. Putting the -1 in the \[y _{1}\] spot allows it to become positive because two negatives make a positive.

OpenStudy (anonymous):

So, (0, -1) and (2, 3) would look like this with their appropriate values: (x, -y) and (x, y) right?

OpenStudy (anonymous):

But you will get the same answer no matter which set of points you plug first, but always be sure to keep the order and correct to your above question

OpenStudy (anonymous):

so go ahead and solve for the slope (m) and tell me what you get

OpenStudy (anonymous):

But, the example of the order of my values(in my last comment) are correct, right?

OpenStudy (anonymous):

yes you are correct = )

OpenStudy (anonymous):

Ah! Thank you. I'll try to solve the problem. One minute (:

OpenStudy (anonymous):

now that is only part of it, there is another equation to arrive to the final equation

OpenStudy (anonymous):

once you find the slope I can tell you the next part = )

OpenStudy (anonymous):

And I would multiply \[x{2}\] and \[y{2}\] like any other number with power, right? i.e. 2 x 2 and 3 x 3

OpenStudy (anonymous):

Oh I'm sorry, the subscript 1 and 2 are only there to decipher between the different values those aren't exponents

OpenStudy (anonymous):

Ohhh ok, thank you. I couldn't remember the right terminology either, sorry. But, I think it's: \[\frac{ 4 }{ -1 }\]

OpenStudy (anonymous):

here let me do this \[\frac{ y-y }{ x-x }\] do you see how this can look like the answer would be 0 right? that's why they put the 1 and 2 on the x and y to determine between the two sets of points

OpenStudy (anonymous):

your numerator is correct double check your denominator (2-0) = ?

OpenStudy (anonymous):

Pff... Sorry. I mixed my numbers up >.<

OpenStudy (anonymous):

no apologies = ) I was just on here earlier getting help and I did the same thing lol ; D

OpenStudy (anonymous):

Essentially it comes out to \[\frac{ 2 }{ 1 }\]

OpenStudy (anonymous):

lol I guess we all make mistakes haha! It's how we learn.

OpenStudy (anonymous):

correct = ) so that is your "m" or slope

OpenStudy (anonymous):

yay! Alright, so that should be written out like so: \[y =2x +b\]

OpenStudy (anonymous):

And b is the 'base', right?

OpenStudy (anonymous):

the next equation you will need is: \[y-y _{1} = m (x-x _{1})\] Now don't get confused that the \[y _{1}\] and \[x _{1}\] are the same as the ones above you can choose either set of points to plug in. The regular y and x are the ones that will remain in the equation after plugging in the set of points which will give you the y=mx+b. Does this make sense so far?

OpenStudy (anonymous):

I can't remember if it is called the base but that will be your y-intercept once you graph it

OpenStudy (anonymous):

Ok so i chose the points (2,3) to use bc I like positive numbers lol Plug in the set of points and what you got for m into the equation \[y-(3)=(2)(x-(2))\] Tell me what you get after solving that

OpenStudy (anonymous):

@masukasu

OpenStudy (anonymous):

Sorry. Someone had to use my laptop.

OpenStudy (anonymous):

that's ok = )

OpenStudy (anonymous):

Ok, I'm going to solve that. One minute (:

OpenStudy (anonymous):

Wait... Do I NEED the parenthesis around the numbers? Or does it not really change anything?

OpenStudy (anonymous):

It's just easier to see if you have any sign changes. Now say you used the other set of points instead of the positive numbers you would end up with this: \[y-(-1) = (2)(x-0) \] It's safe to have that parentheses around the -1 bc the sign of it changes to y+1 which will screw things up when you have to solve for y

OpenStudy (anonymous):

but it will only screw things up when you have to solve for y if you forget to change the sign of the -1.

OpenStudy (anonymous):

This equation is a little confusing. I'm used to it looking like \[2\left( x -2 \right)=y -\left( 3 \right)\] but even then, it's a little difficult.

OpenStudy (anonymous):

ok so the only difference is the sides of the equation are on opposite sides = ) do you know how to factor? I'll help you step by step = )

OpenStudy (anonymous):

I would bring the 3 over and it would become a +3, then it would look like this:\[2\left( x-2 \right)+3=y\] I think

OpenStudy (anonymous):

Factoring... Uh lol Is that factoring? I can't remember. Am I able to reverse the equation like I did though?

OpenStudy (anonymous):

that is a correct move but that's not factoring. Take a look at the 2(x-2): when you see something like this you use factoring where you take the front number and multiply it by both numbers inside the parentheses. So you would have 2*x - 4 then tack on what's left in the equation: 2x-4+3=y Then simplify

OpenStudy (anonymous):

OH! I remember now!

OpenStudy (anonymous):

Well... that part at least

OpenStudy (anonymous):